For this problem I am to solve for the specified variable. I am not even sure how to start. Can you please help. Thank you so much!!!
x=vt- 1/2at^2 ; t
Thank you again!!

Answer by stanbon(75887)

x=vt- 1/2at^2
(1/2)t^2-vt+x = 0
Use the Quadratic formula:
t = [v +- sqrt(v^2 -4*(1/2)x)]/(2*1/2)
t = [v +- sqrt(v^2 - 2x)]
t = v + sqrt(v^2-2x) or t = v- sqrt(v^2-2x)
===============
Cheers,
Stan H.